OpenStudy (m_m13):

will medal! Solve the equation for y. 4x – 5y = 9

OpenStudy (photon336):

Okay, first step we need to move the x terms to one side

OpenStudy (m_m13):

ok

OpenStudy (photon336):

4x-5y = 9 To get this below: \[5y =(4x-9)\] now we divide both sides by 5 \[y = \frac{ 4 }{ 5 }x- \frac{ 9 }{ 5}\]

OpenStudy (mrnood):

HMMM to solve fo y you need to move the Y terms to one side, and all the rest on the other

OpenStudy (m_m13):

ok